# Descend or Rewind? Stochastic Gradient Descent Unlearning

### v1

@@ B.6 SGD-R2D
To prove Theorem 3, we need to determine $\Sigma$ such that $\mathbb{E}[\lVert\theta^{\prime}_{t}-\theta^{\prime\prime}_{K}\rVert]\leq\Sigma$. This result can be combined with the relaxed Gaussian mechanism (Lemma 12) to yield the $(\epsilon,\delta)$ guarantee in Theorem 3. We can determine $\Sigma$ for contracting, semi-contracting, and expanding systems, which correspond to strongly convex, convex, and nonconvex loss functions (Lemma 15). This yields our final noise guarantees in Theorem 3.
As discussed in the main body, we require Lemma 19, which bounds the unlearning bias in terms of the second moment of the stochastic noise.
Lemma 19.
@@
Proof. See Appendix B.6.1.
In addition, we require Lemma 20, a classic result which states that assuming the noise is relatively bounded (Assumption 4), the gradient norms of the unbiased SGD iterates are bounded. Lemma 19 and 20 allow us to isolate and capture the accumulated disturbances from noise and bias, which are upper bounded by a linear term in $T$ and the loss at initialization. This can be combined with Assumption 5 to achieve a dataset-independent bound.
Lemma 20.
(Convergence of SGD) Let $\theta_{t}$ represent SGD iterates (2) on a loss function $\mathcal{L}_{\mathcal{D}}$. Suppose Assumptions 1 and 4 are satisfied. Then for $\theta_{t+1}=\theta_{t}-\eta g_{\mathcal{B}_{t}}(\theta_{t})$ and $\eta\leq\frac{1}{LB}$, we have
@@
Proof. See Appendix B.6.2.
We use the above results to prove the following Lemma 21 that bounds the expected distance between $\theta^{\prime}_{t}$ and $\theta_{t}$.
Lemma 21.
Suppose that the loss function $\ell$ satisfies Assumptions 1, 4, and 5. Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the following property for some $\gamma>0$,
@@
Proof. See Appendix B.6.3.
Lemma 21 allows us to prove Theorem 22, which can be combined with Lemma 15 and Lemma 12 to yield the result in Theorem 3.
Theorem 22.
Suppose that the loss function $\ell$ satisfies Assumptions 1, 4, and 5. Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the property that for some $\gamma>0$ $\lVert T^{z}_{\eta}(\theta)-T^{z}_{\eta}(\theta^{\prime})\rVert\leq\gamma\lVert\theta-\theta^{\prime}\rVert$. Let $\theta_{t}$, $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$ denote the training, retraining, and unlearning iterates as defined in (22), (23) and (24) respectively. If $\gamma\neq 1$, there exists a coupling of $\theta_{t}$, $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$ such that
@@
Lemma 21 shows that the expected distance between $\theta_{t}$, $\theta^{\prime}_{t}$, is bounded as a function of $t$. So for $\gamma\neq 1$, we have
@@
From Lemma 16, we have that we can choose a coupling of $\{\theta^{\prime}_{t}\}_{t=T-K}^{T}$, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$, such that
@@
B.6.1 Proof of Lemma 19
To bound the unlearning bias, we first require the following result from [Polyanskiy_Wu_2025] (Theorem 7.26 and Example 7.4), which follows from the Radon-Nikodym theorem and Cauchy-Schwarz inequality.
Lemma 23.
@@ B.6.1 Proof of Lemma 19
Now we can proceed with the proof of Lemma 19.
@@
In the following, we provide an explicit proof that does not require any prior knowledge for easy understanding. However, we note that the result ultimately follows from defining the empirical distributions of $\mathcal{D}$ and $\mathcal{D}^{\prime}$ as $P_{\mathcal{D}}$ and $P_{\mathcal{D}^{\prime}}$, and applying Lemma 23.
@@
The theory underlying Lemma 19 is that we use the second moment bound to write the difference in terms of the $\chi^{2}$-divergence between $P_{\mathcal{D}}$ and $P_{\mathcal{D}^{\prime}}$, where $w(x)$ represents a discrete version of the Radon Nikodym derivative $\frac{dP_{\mathcal{D}^{\prime}}}{dP_{\mathcal{D}}}$. The $\chi^{2}$-divergence then evaluates to $\frac{m}{n-m}$.
B.6.2 Proof of Lemma 20
@@
By Lipschitz smoothness (Lemma 8), we have
@@
By Assumption 4 and $\eta\leq\frac{1}{LB}$, we have
@@ B.6.3 Proof of Lemma 21
where (2) represents the unlearning bias and (3) represents the noise. By (26) and (16) of Lemma 15, we have
@@
By our bias bound (Lemma 19) and relative noise bound assumption (Assumption 4), we have
@@
We can also use Assumption 4 to bound the noise terms in $(3)$ as follows,
@@
Combining the above bounds and utilizing the fact that $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and Lemma 6, we have
@@
B.7 SGD-D2D
In this section, we prove Theorem 4, which establishes certified unlearning for SGD-D2D on strongly convex functions. We first establish preliminaries and provide an overall proof sketch. In Appendix B.7.1 we prove Lemma 24, and in Appendix B.7.2 we prove Lemma 25, which are necessary for achieving the final result. Finally, in Appendix B.7.3, we finish the proof of Theorem 4.
@@ B.7 SGD-D2D
The first step of the proof is to show that during training the biased SGD iterates $\{\theta_{t}\}_{t=0}^{T}$ will contract to be within some neighborhood of $\theta^{*^{\prime}}$, the optimum of $\mathcal{L}_{\mathcal{D}^{\prime}}$ (Lemma 25). This result replicates the general result in [AjalloeianBiasedSGD] showing linear convergence of biased SGD as long as the following conditions hold: i) the loss function is PL and Lipschitz smooth, ii) the noise satisfies a (relative) bound as in Assumption 4, and iii) there exists constants $D\geq 0$, $0\leq M<1$, such that the bias at time step $t$, denoted as $d_{t}$, is relatively bounded as follows,
@@
We show that we satisfy the bias bound condition (31) if the proportion of unlearned data is small enough in Lemma 24.
Lemma 24.
Suppose that Assumption 4 holds and $\frac{m}{n}\leq\frac{1}{6B+1}$. Then for all $\theta\in\mathbb{R}^{d}$, the unlearning bias is bounded as follows,
@@
Proof. See Appendix B.7.1.
Lemma 25.
@@
Proof. See Appendix B.7.2.
The second step is to show that the unbiased SGD iterates $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ also converge close to $\theta^{*^{\prime}}$ through the classic convergence analysis. Finally, the last step is to show that during unlearning, the unbiased SGD iterates $\{\theta_{t}^{\prime\prime}\}_{t=0}^{K}$ will converge closer $\theta^{*^{\prime}}$, closing the gap up to some neighborhood determined by the stochasticity. By tracking the progress of the loss value $\mathcal{L}_{\mathcal{D}^{\prime}}$ and leveraging the quadratic growth condition (10), we achieve a second-moment bound on $\lVert\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}\rVert$, which can be combined with Lemma 12 to yield $(\epsilon,\delta)$-indistinguishability. In Appendix B.7.3, we carry out these proof components to achieve the result in Theorem 4. This approach is unique to strongly convex functions and cannot be achieved for convex and nonconvex functions.
B.7.1 Proof of Lemma 24
@@
Combining Lemma 19 with Assumption 4, we have
@@ B.7.2 Proof of Lemma 25
By Lipschitz smoothness (Lemma 8), we have
@@
By the fact that $\lVert a+b\rVert^{2}-2a^{T}(a+b)=-\lVert a\rVert^{2}+\lVert b\rVert^{2}$, the bias cross term $\nabla\mathcal{L}_{\mathcal{D}^{\prime}}(\theta_{t})^{T}d_{t}$ can get “folded” into the norm squared of $d_{t}$, which is bounded away with Lemma 24. We have
@@
Assuming that $\frac{m}{n}\leq\frac{1}{1+6B}$, we can bound the bias term by Lemma 24 as follows,
@@
We can use the Polyak–Łojasiewicz (PL) property of strongly convex functions (Assumption 3 and Definition 4) to bound in terms of $\mathcal{L}^{*}_{\mathcal{D}^{\prime}}=\inf_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{\mathcal{D}^{\prime}}(\theta)$.
@@ B.7.3 Proof of Theorem 4
Now we analyze the linear convergence of $\theta^{\prime\prime}_{t}$ on $\mathcal{L}_{\mathcal{D}^{\prime}}$, leading to a similar result depending on $\mathbb{E}[\mathcal{L}_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{0})]-\mathcal{L}_{\mathcal{D}^{\prime}}^{*}$. This allows us to plug in the results from Lemma 25. We have
@@
By quadratic growth (Lemma 10), we have
@@
Then by Lemma 12, we obtain $(\epsilon,2\delta)$-certified unlearning if we add Gaussian noise with standard deviation
$\sigma=\frac{\Sigma}{\epsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}}.$
Appendix C Implementation
Theorem 26.
[khaled2020bettertheorysgdnonconvex] [sebbouh21a] Suppose that Assumption 1 holds and the loss function $\mathcal{L}_{\mathcal{D}}$ has the finite-sum structure
$\mathcal{L}_{\mathcal{D}}(\theta)=\frac{1}{n}\sum_{z\in\mathcal{D}}\ell(z;\theta),$
and let $\mathcal{L}_{\mathcal{D}}^{*}$ represent its minimum value. Then for $\Delta^{inf}_{\mathcal{D}}=\frac{1}{n}\sum_{i=1}^{n}(\mathcal{L}_{\mathcal{D}}^{*}-\ell_{i}^{*})$
$\mathbb{E}_{\mathcal{B}\sim\mathcal{D}}[\lVert g_{\mathcal{B}}(\theta)\rVert^{2}]\leq A(\mathcal{L}_{\mathcal{D}}(\theta)-\mathcal{L}_{\mathcal{D}}^{*})+B\lVert\nabla\mathcal{L}_{\mathcal{D}}(\theta)\rVert^{2}+C$ (33)
for all $\theta\in\Theta$, where $|\mathcal{B}|=b$, and $A$, $B$, and $C$ are constants determined by the batch sampling scheme:
  • With replacement: $A=\frac{L}{b}$, $B=1-\frac{1}{b}$, $C=2A\Delta^{inf}_{\mathcal{D}}$,
  • Without replacement: $A=\frac{n-b}{b(n-1)}L$, $B=\frac{n(b-1)}{b(n-1)}$, $C=2A\Delta^{inf}_{\mathcal{D}}$,
and the analogous results hold for $\mathcal{D}^{\prime}$ and $\mathcal{L}_{\mathcal{D}^{\prime}}$.
Theorem 26 follows from Proposition 3 in [khaled2020bettertheorysgdnonconvex]. It states that we can bound the second moment of the stochastic gradient estimator due to the implications of the finite-sum loss function structure and Assumption 1. In addition, it states that the constants $A$ and $B$ decrease and increase respectively with increasing batch size $b$. This condition was first introduced in [khaled2020bettertheorysgdnonconvex] as the expected smoothness assumption, more recently referred to as the ABC assumption [sebbouh21a][demidovich2023guidezoobiasedsgd][gower2021sgdstructurednonconvexfunctions][garrigos2024handbookconvergencetheoremsstochastic]. We note that our analysis may be extended to other sampling schemes covered by [khaled2020bettertheorysgdnonconvex], such as weighted sampling.
Theorem 26 states that the gradient noise bound depends on $\Delta^{inf}_{\mathcal{D}}$, which we refer to as the interpolation constant because it is zero when the model can perfectly interpolate between all data samples.

### v2

@@ A.6 SGD-R2D
To prove Theorem 4.9, we need to determine $\Sigma$ such that $\mathbb{E}[\lVert\theta^{\prime}_{t}-\theta^{\prime\prime}_{K}\rVert]\leq\Sigma$. This result can be combined with Lemma 4.4 to yield the $(\varepsilon,\delta)$ guarantee in Theorem 4.9. We can determine $\Sigma$ for contracting, semi-contracting, and expanding systems, which correspond to strongly convex, convex, and nonconvex loss functions (Lemma A.11). This yields our final noise guarantees in Theorem 4.9.
As discussed in the main body, we require Lemma A.15, which bounds the unlearning bias in terms of the second moment of the stochastic noise.
Lemma A.15.
@@
Proof. See Appendix A.6.1.
In addition, we require Lemma A.16, a classic result which states that assuming the noise is relatively bounded (Assumption 4.7), the gradient norms of the unbiased SGD iterates are bounded. Lemma A.15 and A.16 allow us to isolate and capture the accumulated disturbances from noise and bias, which are upper bounded by a linear term in $T$ and the loss at initialization. This can be combined with Assumption 4.8 to achieve a dataset-independent bound.
Lemma A.16.
(Convergence of SGD) Let $\theta_{t}$ represent SGD iterates (2) on a loss function $\mathcal{L}_{\mathcal{D}}$. Suppose Assumptions 4.1 and 4.7 are satisfied. Then for $\theta_{t+1}=\theta_{t}-\eta g_{\mathcal{B}_{t}}(\theta_{t})$ and $\eta\leq\frac{1}{LB}$, we have
@@
Proof. See Appendix A.6.2.
We use the above results to prove the following Lemma A.17 that bounds the expected distance between $\theta^{\prime}_{t}$ and $\theta_{t}$.
Lemma A.17.
Suppose that the loss function $\ell$ satisfies Assumptions 4.1, 4.7, and 4.8. Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the following property for some $\gamma>0$,
@@
Proof. See Appendix A.6.3.
Lemma A.17 allows us to prove Theorem A.18, which can be combined with Lemma A.11 and Lemma 4.4 to yield the result in Theorem 4.9.
Theorem A.18.
Suppose that the loss function $\ell$ satisfies Assumptions 4.1, 4.7, and 4.8. Moreover, suppose for all $z\in\mathcal{Z}$ and for any $\theta,\theta^{\prime}\in\mathbb{R}^{d}$, the gradient descent map $T^{z}_{\eta}(\theta)=\theta-\eta\nabla\ell(z;\theta)$ satisfies the property that for some $\gamma>0$ $\lVert T^{z}_{\eta}(\theta)-T^{z}_{\eta}(\theta^{\prime})\rVert\leq\gamma\lVert\theta-\theta^{\prime}\rVert$. Let $\theta_{t}$, $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$ denote the training, retraining, and unlearning iterates as defined in (20), (21) and (22) respectively. If $\gamma\neq 1$, there exists a coupling of $\theta_{t}$, $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$ such that
@@
Lemma A.17 shows that the expected distance between $\theta_{t}$, $\theta^{\prime}_{t}$, is bounded as a function of $t$. So for $\gamma\neq 1$, we have
@@
From Lemma A.12, we have that we can choose a coupling of $\{\theta^{\prime}_{t}\}_{t=T-K}^{T}$, $\{\theta^{\prime\prime}_{t}\}_{t=0}^{K}$, such that
@@
A.6.1 Proof of Lemma A.15
To bound the unlearning bias, we first require the following result from (Polyanskiy and Wu, 2025) (Theorem 7.26 and Example 7.4), which follows from the Radon-Nikodym theorem and Cauchy-Schwarz inequality.
Lemma A.19.
@@
Now we can proceed with the proof of Lemma A.15.
@@
In the following, we provide an explicit proof that does not require any prior knowledge for easy understanding. However, we note that the result ultimately follows from defining the empirical distributions of $\mathcal{D}$ and $\mathcal{D}^{\prime}$ as $P_{\mathcal{D}}$ and $P_{\mathcal{D}^{\prime}}$, and applying Lemma A.19.
@@
The theory underlying Lemma A.15 is that we use the second moment bound to write the difference in terms of the $\chi^{2}$-divergence between $P_{\mathcal{D}}$ and $P_{\mathcal{D}^{\prime}}$, where $w(x)$ represents a discrete version of the Radon Nikodym derivative $\frac{dP_{\mathcal{D}^{\prime}}}{dP_{\mathcal{D}}}$. The $\chi^{2}$-divergence then evaluates to $\frac{m}{n-m}$.
A.6.2 Proof of Lemma A.16
@@
By Lipschitz smoothness (Lemma A.4), we have
@@
By Assumption 4.7 and $\eta\leq\frac{1}{LB}$, we have
@@
where (2) represents the unlearning bias and (3) represents the noise. By (24) and (14) of Lemma A.11, we have
@@
By our bias bound (Lemma A.15) and relative noise bound assumption (Assumption 4.7), we have
@@
We can also use Assumption 4.7 to bound the noise terms in $(3)$ as follows,
@@
Combining the above bounds and utilizing the fact that $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and Lemma A.2, we have
@@
A.7 SGD-D2D
In this section, we prove Theorem 4.10, which establishes certified unlearning for SGD-D2D on strongly convex functions. We first establish preliminaries and provide an overall proof sketch. In Appendix A.7.1 we prove Lemma A.20, and in Appendix A.7.2 we prove Lemma A.21, which are necessary for achieving the final result. Finally, in Appendix A.7.3, we finish the proof of Theorem 4.10.
@@ A.7 SGD-D2D
The first step of the proof is to show that during training the biased SGD iterates $\{\theta_{t}\}_{t=0}^{T}$ will contract to be within some neighborhood of $\theta^{*^{\prime}}$, the optimum of $\mathcal{L}_{\mathcal{D}^{\prime}}$ (Lemma A.21). This result replicates the general result in (Ajalloeian and Stich, 2020) showing linear convergence of biased SGD as long as the following conditions hold: i) the loss function is PL and Lipschitz smooth, ii) the noise satisfies a (relative) bound as in Assumption 4.7, and iii) there exists constants $D\geq 0$, $0\leq M<1$, such that the bias at time step $t$, denoted as $d_{t}$, is relatively bounded as follows,
@@
We show that we satisfy the bias bound condition (29) if the proportion of unlearned data is small enough in Lemma A.20.
Lemma A.20.
Suppose that Assumption 4.7 holds and $\frac{m}{n}\leq\frac{1}{6B+1}$. Then for all $\theta\in\mathbb{R}^{d}$, the unlearning bias is bounded as follows,
@@
Proof. See Appendix A.7.1.
Lemma A.21.
@@
Proof. See Appendix A.7.2.
The second step is to show that the unbiased SGD iterates $\{\theta^{\prime}_{t}\}_{t=0}^{T}$ also converge close to $\theta^{*^{\prime}}$ through the classic convergence analysis. Finally, the last step is to show that during unlearning, the unbiased SGD iterates $\{\theta_{t}^{\prime\prime}\}_{t=0}^{K}$ will converge closer $\theta^{*^{\prime}}$, closing the gap up to some neighborhood determined by the stochasticity. By tracking the progress of the loss value $\mathcal{L}_{\mathcal{D}^{\prime}}$ and leveraging the quadratic growth condition (10), we achieve a second-moment bound on $\lVert\theta^{\prime}_{T}-\theta^{\prime\prime}_{K}\rVert$, which can be combined with Lemma 4.4 to yield $(\varepsilon,\delta)$-indistinguishability. In Appendix A.7.3, we carry out these proof components to achieve the result in Theorem 4.10. This approach is unique to strongly convex functions and cannot be achieved for convex and nonconvex functions.
A.7.1 Proof of Lemma A.20
@@
Combining Lemma A.15 with Assumption 4.7, we have
@@
By Lipschitz smoothness (Lemma A.4), we have
@@
By the fact that $\lVert a+b\rVert^{2}-2a^{T}(a+b)=-\lVert a\rVert^{2}+\lVert b\rVert^{2}$, the bias cross term $\nabla\mathcal{L}_{\mathcal{D}^{\prime}}(\theta_{t})^{T}d_{t}$ can get “folded” into the norm squared of $d_{t}$, which is bounded away with Lemma A.20. We have
@@
Assuming that $\frac{m}{n}\leq\frac{1}{1+6B}$, we can bound the bias term by Lemma A.20 as follows,
@@
We can use the Polyak–Łojasiewicz (PL) property of strongly convex functions (Assumption 4.3 and Definition A.6) to bound in terms of $\mathcal{L}^{*}_{\mathcal{D}^{\prime}}=\inf_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{\mathcal{D}^{\prime}}(\theta)$.
@@
Now we analyze the linear convergence of $\theta^{\prime\prime}_{t}$ on $\mathcal{L}_{\mathcal{D}^{\prime}}$, leading to a similar result depending on $\mathbb{E}[\mathcal{L}_{\mathcal{D}^{\prime}}(\theta^{\prime\prime}_{0})]-\mathcal{L}_{\mathcal{D}^{\prime}}^{*}$. This allows us to plug in the results from Lemma A.21. We have
@@
By quadratic growth (Lemma A.7), we have
@@
Then by Lemma 4.4, we obtain $(\varepsilon,2\delta)$-certified unlearning if we add Gaussian noise with standard deviation
$\sigma=\frac{\Sigma}{\varepsilon}\sqrt{\frac{2\log(1.25/\delta)}{\delta}}.$
Appendix B Experiments
B.1 Implementation Details
We follow the experimental setup for unlearning detailed in Section 4.1 and Appendix B of (Mu and Klabjan, 2025), including dataset preparation, model architecture, unlearning procedure, and MIA implementations. We implement PSGD by projecting iterates onto a ball of radius $R$ centered on the origin. We use the hyperparameters listed in Table 2. Code is open-sourced at the anonymous GitHub repository https://anonymous.4open.science/r/r2d2-3753/.
All experiments were run using PyTorch 2.5.0 and CUDA 12.1, on an Intel(R) Core(TM) i7-6850K CPU (3.60GHz) with an NVIDIA GeForce GTX 1080 GPU (8 GB VRAM) or on an Intel(R) Xeon(R) Silver 4208 CPU (2.10GHz) with an NVIDIA RTX A6000 GPU (48 GB).
Table 2: R2D Experiment parameters for the eICU and Lacuna-100 datasets.
Parameter eICU and MLP Lacuna-100 and ResNet-18
Size of training dataset $n$ 94449 32000
Number of users 119282 100
Percent data unlearned $\sim$ 1% $\sim$ 2%
Number of model parameters $d$ 136386 11160258
Batch size 64 64
$L$ 0.059955
$G$ 0.820322
$\eta$ 0.001 0.01
Number of training epochs 48 43
$R$ 10 50
B.2 Privacy-utility-complexity tradeoff
To demonstrate the real-world tradeoffs implied by the derived guarantees in Theorem 4.9, we vary $\varepsilon$, compute the required $\sigma$, and examine the model performance and MIA success. We estimate the constants $G$ and $L$ by sampling uniformly from the parameter space with radius $R$. Figure 3 demonstrates the privacy-utility-complexity tradeoff of R2D on the nonconvex loss function on the eICU dataset, where “Rewind Percent” is computed as $\frac{K}{T}\%$. As expected, the behavior trends in accordance with the derived relationship between $\varepsilon$, $K$, and $\sigma$.
B.3 Baseline Comparison
We are also interested in comparing R2D with the existing Privacy Amplification By Iteration PABI) algorithm (Koloskova et al., 2025) because of its nonconvex unlearning guarantees. We do not compare against Langevin Unlearning (Chien et al., 2024b) because, as stated in their work, “the non-convex unlearning bound… currently is not tight enough to be applied in practice due to its exponential dependence on various hyperparameters,” including the projection radius.
PABI follows a D2D-like structure but involves clipping and added noise at every step. One advantage of this approach is that it leverages amplification to stop noise injections early, as long as the privacy budget is achieved. After this point the algorithm continues with noiseless finetuning steps to improve accuracy.
We implement PABI and R2D with $\varepsilon=1e7$ and $\delta=0.2$ and the same learning rate, batch size, projection radius, and number of iterations, and we compare their unlearning performance. We use a clipping radius of $1$ and $0.01$ weight decay.
Table 3 displays the results of our experiments. We observe that PABI achieves a better model performance due to a lower noise requirement and the noiseless finetuning steps. However, as with D2D, these finetuning steps also improve the performance on $\mathcal{D}_{unlearn}$, which may even increase after unlearning. In contrast, the noise for R2D is larger and degrades model performance, but it also reduces the gap between the model performance on $\mathcal{D}_{unlearn}$ and the performance on the other sets. As for the membership inference attacks, R2D defends more successfully. However, it is worth noting that MIAs are a somewhat flawed metric for unlearning and their success correlates with model performance.
Figure 3: Privacy-utility-complexity tradeoff of PSGD R2D $(\varepsilon,\delta)$-unlearning on the eICU dataset. Rewind Percent is computed as $\frac{K}{T}\times 100\%$, where $K$ is the number of unlearning iterations and $T$ is the total number of training iterations.
Table 3: Comparison of Certified Unlearning Algorithms on the eICU Dataset
Model AUC MIA AUC
Algorithm $\mathcal{D}_{ood}$ $\mathcal{D}_{retain}$ $\mathcal{D}_{unlearn}$ $\mathcal{D}_{test}$ Classic MIA MIA-U
Original Model
0.728316 0.741235 0.753042 0.739546
Certified Unlearning Algorithms
[…]
PABI (35%) 0.728660 0.741511 0.752976 0.739329 0.506902 $\pm$0.028156 0.5125 $\pm$0.025596
[…]
B.4 Hyperparameters $R$ and $b$
We conduct additional experiments to ensure appropriate choices of projection radius $R$ and batch size $b$. For $R$, following the precedent in (Zhang et al., 2024), we implement projected SGD on a ball of radius $R$ to maintain the strictness of our results while choosing $R$ large enough to minimize impact on model utility. We assess this by examining the training loss and error for varying choices of $R$. For these experiments, we train the model (with batch size $b=64$) and perform model selection of the parameters with lowest validation loss. As shown in Table 4, the choices of $R=10$ and $R=50$ for eICU and Lacuna-100 are have minimal impact on model performance.
For the batch size $b$, we desire a batch size that is small enough to highlight the effects of stochasticity but large enough so that training is stable enough to yield good model performance. We consider the training and test loss for varying choices of $b$, where for each the model is trained with the same number of iterations. Tables 5 and 6 demonstrate the impact of varying batch sizes on the performance on the validation and training sets.
Table 4: Choice of projection radius $R$ vs. model performance for eICU (left) and Lacuna-100 (right).
eICU Dataset
$R$ Train Error Train Loss $G$ $L$
1 0.403646 0.679236 0.722412 0.077918
2 0.394975 0.647507 0.726842 0.076906
5 0.313513 0.588481 0.745951 0.071007
10 0.308939 0.581918 0.820322 0.059955
15 0.308907 0.581422 1.084669 0.057535
20 0.308907 0.581422 1.919703 0.082583
Lacuna-100 Dataset
$R$ Train Error Train Loss
20 0.502000 0.693194
30 0.059719 0.148735
40 0.003563 0.013695
50 0.020094 0.058069
60 0.012969 0.045947
70 0.012969 0.045947
Table 5: Choice of batch size $b$ vs. model performance for the eICU dataset.
Batch Size Train Loss Train Error Validation Loss Validation Error
5 8 0.416120 0.250000 0.583045 0.308373
4 16 0.513511 0.250000 0.582881 0.308373
3 32 0.493724 0.250000 0.581867 0.307314
2 64 0.551778 0.296875 0.581444 0.306467
1 128 0.557833 0.265625 0.581424 0.306425
0 256 0.588656 0.313111 0.588421 0.309813
Table 6: Choice of batch size $b$ vs. model performance for the Lacuna-100 dataset.
Batch Size Train Loss Train Error Validation Loss Validation Error
5 8 0.315790 0.250000 0.244619 0.096125
4 16 0.098530 0.000000 0.184725 0.074375
3 32 0.020336 0.000000 0.201231 0.056375
2 64 0.010567 0.000000 0.220710 0.049125
1 128 0.000255 0.000000 0.342793 0.051500
0 256 0.126300 0.048500 0.230133 0.093125
Appendix C Potential Future Directions
There are several potential future directions of this work. First, we note that other certified unlearning algorithms, such as (Chien et al., 2024a) and (Koloskova et al., 2025), follow a D2D-like framework with clipped or noisy gradient updates, instead of vanilla SGD. Combining their approaches with R2D instead could yield improved guarantees, especially on nonconvex functions. Second, our results for PSGD rely on sampling with replacement to achieve tighter bounds. Whenever a sample from $Z$ is selected during the generation of $\theta_{t}$, we can replace it with an i.i.d. sample from $\mathcal{D}^{\prime}$ for producing $\theta^{\prime}_{t}$, maintaining a “valid” run of $\theta^{\prime}_{t}$. However, if sampling occurs without replacement, then drawing one sample affects the probability of the next. Developing the corresponding theory for sampling without replacement, which is more commonly implemented in practice, would be a useful future direction.
Appendix D Additional Discussion of Related Work
Differential Privacy.
The concept of certified unlearning is directly built on the existing mathematical framework of differential privacy (DP) (Dwork et al., 2006). DP formalizes privacy guarantees by leveraging noise injection to limit the impact of the inclusion or exclusion of any one data sample on the output of the algorithm. The level of privacy is controlled by the parameters $\varepsilon$ and $\delta$, as stated in the definition below.
Definition D.1.
(Dwork and Roth, 2014) A randomized algorithm $\mathcal{M}$ with domain $\mathbb{N}^{|\mathcal{X}|}$ is $(\varepsilon,\delta)$-differentially private if for all $S\subset Range(\mathcal{M})$ and for all adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}\in\mathbb{N}^{|\mathcal{X}|}$
$\mathbb{P}[\mathcal{M}(\mathcal{D})\in S]\leq e^{\varepsilon}\mathbb{P}[\mathcal{M}(\mathcal{D}^{\prime})\in S]+\delta.$
A common DP technique is the Gaussian mechanism, which scales Gaussian noise to the deterministic global sensitivity, a worst-case bound on how much the algorithm output shifts when the underlying dataset is changed (Dwork and Roth, 2014). This technique has been applied to privatizing deterministic functions such as dataset queries or full-batch gradient descent, as well as randomized algorithms such as SGD (Abadi et al., 2016; Wu et al., 2017; Zhang et al., 2017). In particular, to bound the sensitivity, the well-known DP-SGD algorithm (Abadi et al., 2016) clips the gradient at each step to minimize the impact of any single data sample. Alternatively, the black-box SGD algorithm (Wu et al., 2017) requires a uniformly bounded gradient and strongly convex loss function.
Biased SGD.
Our work relies on interpreting the certified unlearning problem as a form of biased or disturbed gradient descent. There are two complementary viewpoints in the literature. First, several works analyze biased SGD from a classic optimization perspective to achieve convergence to a minimum or stationary point. Typically, these works establish minimal viable assumptions on the bias or noise to achieve convergence. For example, (Demidovich et al., 2023) summarizes biased SGD results for nonconvex and strongly convex (or PL) loss functions. In addition, (Hu et al., 2016) considers bandit convex optimization with absolutely bounded bias and on a bounded convex domain, and (Driggs et al., 2022) considers biased SGD schemes that satisfy a specific “memory-biased” property, like SARAH and SVRG. In our work, we utilize the groundwork laid in (Ajalloeian and Stich, 2020) to show unlearning for D2D on strongly convex functions.
The second viewpoint of biased SGD is through the lens of nonlinear contraction theory (Kozachkov et al., 2022) (Sontag, 2021), characterizing the bias at every step as disturbances to a noisy gradient system. The trajectory stability, which determines how far a gradient trajectory will diverge when disturbed, can be used to bound the distance between SGD training on two different loss functions. We leverage well-known properties of these systems to analyze stochastic R2D unlearning.